OpenStudy (anonymous):

help find Sin(B) ,Cos(B), and Tan(B) for 1.) π/6

OpenStudy (anonymous):

can you simplify your question again?

OpenStudy (mertsj):

Have you studied the unit circle?

OpenStudy (anonymous):

i know how to use the unit circle but it does not make sense when their asking for (B)

OpenStudy (mertsj):

B is pi/6 They are asking for: \[\sin \frac{\pi}{6}\] \[\cos \frac{\pi}{6}\] \[\tan \frac{\pi}{6}\]

OpenStudy (anonymous):

that's all their asking for

OpenStudy (mertsj):

yep

OpenStudy (mertsj):

And they are should be contracted as they're NOT their